General Optic Corporation operates a manufacturing plant in Arizona. Due to a significant decline in demand for the product manufactured at the Arizona site, an impairment test is deemed appropriate. Management has acquired the following information for the assets at the plant:
 





Cost$39,500,000
Accumulated depreciation
14,900,000
General’s estimate of the total cash flows to be generated by selling the products
manufactured at its Arizona plant, not discounted to present value

16,400,000

 
The fair value of the Arizona plant is estimated to be $14,500,000.
 
Required:
1. Determine the amount of impairment loss.
2. If a loss is indicated, prepare the entry to record the loss.
3. & 4. Determine the amount of impairment loss assuming that the estimated undiscounted sum of future cash flows is (3) $15,500,000 instead of $16,400,000 and (4) $24,850,000 instead of $16,400,000.
Answer #1

Req 1)

step 1) 39,500,000-14,900,000= 24,600,000 book value

step 2) 24,600,000-14,500,000= 10,100,000 impairment loss (answer)


Req 2) 

dr. loss on impairment              10,100,000

dr. accumulated depreciation    14,900,000

  cr. plant assets                                           25,000,000

10,100,000+14,900,000= 25,000,000


Req 3)

bv= 24,600,000 since it exceeds 15,500,000 it is a loss

24,600,000-14,500,000(FV)= 10,100,000 impairment loss (answer)


Req 4)

Since 24,850,000 exceeds bv of 24,600,000 there is no loss.

Answer: 0 impairment loss

answered by: anonymous